Grade 12Mechanics

If F =xi +yj acts on a particle.then find the work done by force in displacing the particle from the point A (1,1) to point B( 2,3)

Aqsa Ahmed
10 Years agoGrade 12

1 Answer

Carmel Emmanuel
ApprovedApproved Tutor Answer10 Years ago
FWe know dW=F.ds
DW=(xi+yj).(dx+dy)
\int dW=\int (xi+yj). (dx+dy)
Take coordinates from moving (1,1) to (2,3) as first from (1,1) to (2,1) then from (2,1) to (2,3)
\int dW=\int_{1}^{2}(xi.dx) + \int_{1}^{3}(yj.dy)
(X/2)2]1+ [(y/2)2]13
finally you get 1.5+4.5=6J